OpenStudy (anonymous):

Please explain this simply for me! Medical x-rays, with a wavelength of about 10^(-10) m, can penetrate completely through your skin. a. Ultraviolet rays, which cause sun burn by penetrating only the top level of skin, have a wavelength about 1,000 times the wavelength of an x-ray. Find the wavelength of ultraviolet rays. Show your work.

OpenStudy (anonymous):

There's a second part - b. The wavelengths of visible light are between 3.8 x 10^(-7) m and 7.6 x 10^(-7) m. Are these wavelengths longer or shorter than the ultraviolet rays? Explain

OpenStudy (aum):

"Ultraviolet rays ..... have a wavelength about 1,000 times the wavelength of an x-ray" Wavelength of x-rays is given to you as 10^(-10) m. Therefore, wavelength of ultraviolet rays = 1000 * 10^(-10) = 10^3 * 10^(-10) = ?

OpenStudy (anonymous):

10^-7

OpenStudy (aum):

Yes, Wavelength of ultraviolet rays = 10^(-7) m.

OpenStudy (aum):

But in scientific notation it is written as follows: Wavelength of ultraviolet rays = 1.0 x 10^(-7) m.

OpenStudy (anonymous):

ok thanks what about the second part

OpenStudy (aum):

Compare 1.0 x 10^(-7) m to 3.8 x 10^(-7) m and 7.6 x 10^(-7) m. Is it less, more, or in between?

OpenStudy (anonymous):

more?

OpenStudy (aum):

They all have 10^(-7) and so you simply need to compare 1.0 to 3.8 and 7.6. Is 1.0 less than 3.8? Is 1.0 more than 7.6? Is 1.0 in between 3.8 and 7.6?

OpenStudy (anonymous):

yes yes and no

OpenStudy (aum):

How is 1.0 more than 7.6?

OpenStudy (anonymous):

sorry i though it said less :p itsno

OpenStudy (aum):

1.0 is less than 3.8. So ultraviolet rays have shorter wavelength compared to visible light. But the questions asks: "Are these wavelengths longer or shorter than the ultraviolet rays?" By these wavelengths the question is referring to visible light. Therefore visible light have longer wavelengths than ultraviolet rays.

OpenStudy (aum):

Visible light have longer wavelengths than ultraviolet rays because 3.8 x 10^(-7) m and 7.6 x 10^(-7) m > 1.0 x 10^(-7) m

OpenStudy (anonymous):

thanks!

OpenStudy (aum):

you are welcome.